### v1

@@ B.4 Scoring service
The prompt starts with the question and the definition of the factor as described in Table 1, specifies the input and output format, and ends with the list of posts encoded in JSON objects, one for each line. The posts are grouped with at most ten items per prompt as a tradeoff between making too many API calls and keeping the response time low. This implementation ensures that a feed containing ten political posts results in 8 requests to OpenAI instead of 80.
These requests run in parallel using a custom connection pool that shares HTTP connections and iterates round-robin on multiple API tokens. Each request takes a few seconds to receive a response, but each connection has a maximum waiting time of 8 seconds to ensure a slow API call does not disrupt the user experience. When the timeout is reached, the event is logged, and the factor is set as non-present. To ensure a fast response from OpenAI, we use GPT-3.5 Turbo, which offers a good tradeoff between speed and accuracy. We use a fixed seed for reproducibility purposes.
Since many popular posts appear in multiple feeds, to avoid scoring multiple times the same text, each post score is cached in a backend database and reused when necessary without including the post in the request to OpenAI.
@@
Table 2: Definition of the eight factors used in the GPT prompts.
Appendix C Participant characteristics
Figure 13: Summary of the participants’ recruitment flow.
To enroll in the experiment, participants had to take a screening task by installing a browser extension. The extension opened X, asked participants to log in if necessary, scrolled through their feed automatically, and verified that at least 5% of the posts in their “For You” feed were about politics or social issues. Fig. 15 summarizes the distribution of users by the fraction of political content. Around 73% of the participants qualified and were invited to the full study.
Figure 14 shows the distributions of some of the participants’ demographic information collected in the pre-survey, including race, education level, perceived socioeconomic status, and income.
Figure 14: Distribution of the participants’ demographic characteristics.
Figure 15: Distribution of the fraction of political content in the feed of the participants screened. The green area represents the participants who were qualified to join the study.
C.1 Covariate Balance
Tables 3 and 4 show univariate analyses of covariate balance in the Reduced and Increased Exposure experiments, respectively. Omnibus test of balance using randomization inference are reported in Section 5. We do not observe any evidence of imbalance.
@@ C.1 Covariate Balance
Table 3: Univariate Covariate Balance: Reduced Exposure Experiment
@@
Table 4: Univariate Covariate Balance: Increased Exposure Experiment
Appendix D Adjustment for multiple comparisons
For the two groups (experiments), we independently implement an FDR adjustment [72], grouping the variables as follows:
@@ Appendix D Adjustment for multiple comparisons
We also use correct for multiple comparisons when examining heterogeneous treatment effects.
@@
Appendix E Supplementary Results
E.1 Affective Polarization: Regression Tables
@@
The preregistered regression analysis for the in-feed effect is reported in Table 5, and the regression analyses for the post-experiment effect are reported in Table 6.
@@
Table 7 and Table 8 show the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-survey or only at least one in-feed survey. For both experiments, the treatment effect remains statistically significant. We note that in this sample there is also no evidence of covariate imbalance (reduce, post-experiment: $p=0.86$; reduce, in-feed: $p=0.59$; increase, post-experiment: $p=0.77$; increase, in-feed: $p=0.18$), asymmetric attrition rates (reduce, post-experiment: $p=0.55$; reduce, in-feed: $p=0.28$; increase, post-experiment: $p=0.4$; increase, in-feed: $p=0.56$), or asymmetric attrition patterns (reduce, post-experiment: $p=0.49$; reduce, in-feed: $p=0.69$; increase, post-experiment: $p=0.78$; increase, in-feed: $p=0.98$). Testing procedures are described in Section 5.3.
@@
Recruitment Platform: CloudResearch $-0.58$ $-1.58$
$[-2.69;1.54]$ $[-3.58;0.43]$
@@
Table 5: Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
@@
Recruitment Platform: CloudResearch $-0.76$ $-2.72^{*}$
$[-2.79;1.28]$ $[-5.06;-0.37]$
@@
Table 6: Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
@@
Recruitment Platform: CloudResearch $-0.58$ $-0.82$
$[-2.68;1.51]$ $[-2.84;1.19]$
@@
Table 7: Full population – Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
@@
Recruitment Platform: CloudResearch $-0.79$ $-3.07^{*}$
$[-2.72;1.14]$ $[-5.30;-0.84]$
@@
Table 8: Full population – Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
E.2 Emotions: Regression Tables
The preregistered regression analysis for the in-feed effect in the reduced exposure is reported in Table 9, while for the increased exposure is summarized in Table 10. The post-experiment effect for the reduced exposure experiment is summarized in Table 11 and for the increased exposure in Table 12.
Table 13, Table 14, Table 15, and Table 16 report the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-survey or only at least one in-feed survey. The results are qualitatively similar, the variation in negative emotions is statistically significant in both experiments: increased in the increased exposure experiment and reduced in the reduced exposure experiment.
@@ E.2 Emotions: Regression Tables
Recruitment Platform: CloudResearch $0.56$ $2.42$ $-5.05^{*}$ $-3.09$
$[-2.16;3.27]$ $[-0.41;5.24]$ $[-8.54;-1.55]$ $[-6.25;0.08]$

### v2

@@ Acknowledgments
The prompt starts with the question and the definition of the factor as described in Table Table S3, specifies the input and output format, and ends with the list of posts encoded in JSON objects, one for each line. The posts are grouped into at most ten items per prompt, as a trade-off between making too many API calls and keeping the response time low. This implementation ensures that a feed containing ten political posts results in 8 requests to OpenAI instead of 80.
These requests run in parallel using a custom connection pool that shares HTTP connections and iterates round-robin on multiple API tokens. Each request takes a few seconds to receive a response, but each connection has a maximum waiting time of 8 seconds to ensure a slow API call does not disrupt the user experience. When the timeout is reached, the event is logged, and the factor is set as non-present. To ensure a fast response from OpenAI, we use GPT-3.5 Turbo, which, at the time of the experiment, offered a good trade-off between speed and accuracy. We use a fixed seed for reproducibility. Prior work validated that this approach matches expert annotations  (?).
Since many popular posts appear in multiple feeds, to avoid scoring the same text multiple times, each post score is cached in a backend database and reused when necessary without including the post in the request to OpenAI.
The model we used at the time of our deployment was not multimodal, so we did not process images in the posts. Current models do afford this capability, and we recommend it for future work in this vein.
@@
Table S3: Definition of the eight factors used in the GPT prompts.
S5 Participant Characteristics
To enroll in the experiment, participants had to take a screening task by installing a browser extension. The extension opened X, asked participants to log in if necessary, scrolled through their feed automatically, and verified that at least 5% of the posts in their “For you” feed were about politics or social issues. Fig. S15 summarizes the distribution of users by the fraction of political content. Around 73% of the participants qualified and were invited to the full study.
Figure S14 shows the distributions of some of the participants’ demographic information collected in the pre-experiment survey, including race, education level, perceived socioeconomic status, and income.
Tables S4 and S5 show univariate analyses of covariate balance in the Reduced and Increased Exposure experiments, respectively. Omnibus balance tests using randomization inference are reported in Section S1. We do not observe any evidence of imbalance.
@@
Table S4: Univariate Covariate Balance: Reduced Exposure Experiment
@@
Table S5: Univariate Covariate Balance: Increased Exposure Experiment
Figure S14: Distribution of demographic characteristics among participants who completed the post-experiment survey.
Figure S15: Distribution of the fraction of political content in the feed of the participants screened. The green area represents the participants who were qualified to join the study.
S6 Adjustment for Multiple Comparisons
We perform False Discovery Rate (FDR) adjustment using the method described in  (?), considering each experiment (Reduced and Increased Exposure) separately and grouping the variables as:
@@
We also perform FRD adjustment when examining heterogeneous treatment effects:
@@
S7 Political Attitude Shifts
As preregistered, we report the effect of the intervention on the seven remaining attitudes we used to identify the AAPA posts. We find that Reduced Exposure to AAPA content reduces the Biased Evaluation of Politicized Facts while increasing the exposure increases the Opposition to Bipartisan Cooperation (Fig. S16). However, none of these effects are significant after adjusting for multiple hypothesis testing as preregistered (Appendix S6). We note that our experiments were designed to detect effects on affective polarization (Appendix S1.3), and are likely not as well-powered to detect effects on these outcomes.
Consistent with (?), our work also shows a larger treatment effect on partisan animosity compared to other factors such as support for undemocratic practices and support for partisan violence. These findings also align with recent studies which have observed that affective polarization is not robustly linked to antidemocratic tendencies  (?).
Figure S16: Treatment effects on political attitudes estimated by regressing the participants’ post-experiment survey responses on the treatment indicator, their pre-experiment survey response to the same question, and an indicator of the platform they were recruited on. The p-values are adjusted for multiple hypothesis testing as preregistered.
Figure S17: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the in-feed measurements. The p-values are adjusted for multiple hypothesis testing as preregistered.
Figure S18: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the post-experiment measurements. The p-values are adjusted for multiple hypothesis testing as preregistered.
S8 Heterogeneous Treatment Effects
Fig. S17 and Fig. S18 show the heterogeneous treatment effects on affective polarization in the two experiments. We use the same regression specifications as in the main analyses, but interact the treatment indicator with the moderator and run one regression per moderator. As in the covariate balance and attrition tests, we coarsened the education, income, and SES ladder covariates to avoid small categories. We consider only a subset of covariates and split them into primary and secondary moderators as preregistered. After adjusting the p-values for multiple hypothesis testing, none of the covariates reach statistical significance.
S9 Engagement
We examine three aspects of engagement: (i) activity frequency and duration, i.e., number of sessions and time spent on the platform, (ii) engagement volume, i.e., absolute number of posts viewed, liked, and reposted by the participants, and (iii) engagement rate, i.e., rate at which the participants liked or reposted the posts they viewed.
Activity frequency and duration. We start by analyzing the number of sessions and the amount of time the participants spent on the platform. To track their passive activity on the platform, we configured our browser extension to send a request to our server every minute while the participant’s browser tab with the X page was active. We delineate the end of a session and the beginning of a new one by identifying gaps of over one hour between consecutive requests. To calculate the time a participant spent on the platform per session, we summed the number of requests the browser sent during the session, excluding periods when the participant browsed other websites. To prevent overcounting unusually long sessions—likely caused by participants leaving their browsers open without actively using the platform—we capped all sessions longer than two hours at two hours. Finally, to compute the total time spent per user, we sum the time spent across all of their sessions.
To estimate the effects of our interventions, we adjust for participants’ baseline activity and recruitment platform, similar to our main analysis (Tables S6 and  S7). We report results from linear models but note that using linear models with log-transformed outcomes, Poisson models, or quasi-Poisson models yields similar results. We also note that changing the gap between sessions from one hour to 45 or 90 minutes, or adjusting the maximum session length from two hours to 1.5 or 2.5 hours, did not substantively affect the results reported below.
In the Reduced Exposure experiment, we find that participants in the treatment group had a similar number of sessions per day (estimate: -0.06; 95% CI: [-0.16, 0.04]), but spent -4.71 minutes less per day (95% CI: [-8.22, -1.20]) on the platform compared to those in the control group. In the Increased Exposure experiment, participants in the treatment group had both a similar number of sessions per day (estimate: -0.04; 95% CI: [-0.16, 0.08]) and spent a similar amount of time (estimate: -1.76; 95% CI: [-6.17, 2.66]) as those in the control group.
Figure S19: Treatment effects on participants’ volume and rate of engagement. The points show average marginal effects, and the error bars represent 95% confidence intervals. Effects on engagement volume are estimated using quasi-Poisson regression models, while effects on engagement rates are estimated using Beta regression.
Engagement volume. Next, we analyze the frequency of the participants’ engagement with posts on the platform. As is common in social media data, we find that the distributions of our measures of absolute engagement are highly skewed: most participants view, like, and repost a small number of posts and a small number of participants view, like, and repost many posts. To appropriately model this skewness in the data, we fit quasi-Poisson models using a similar specification as our main analyses, i.e., adjusting for the participant’s volume of the same type of engagement during the baseline period and the platform they were recruited on. We chose a quasi-Poisson regression as it effectively modeled the overdispersion in the data and was less sensitive to outliers than a negative binomial model. While our discussion focuses on the results obtained with quasi-Poisson models, we report the regression tables of the equivalent analyses with a linear model (Tables S8S13). To ease the interpretation of the effects estimated by the quasi-Poisson models, we report average marginal effects. We count a post as viewed if it was in the viewport of the participant’s browser for at least one second.
In the Reduced Exposure experiment, we find that participants in the treated group viewed -169.52 fewer posts (95% CI: [-325.48, -13.56]) over the 7-day intervention period (Figure S19, left panel). They also liked -24.06 fewer posts (95% CI: [-39.53, -8.59]) compared to those in the control condition. Participants in the treatment condition, however, reposted a similar number of posts (estimate: -4.18; 95% CI: [-9.89, 1.53]) as those in the control condition. In the Increased Exposure experiment, we do not find any statistically significant differences in the number of views (estimate: -132.3; 95% CI: [-292.5, 27.94]), likes (estimate: 10.25; 95% CI: [-5.23, 25.74]), or reposts (estimate: 3.38; 95% CI: [-0.06, 6.82]) between the participants in the treatment and control conditions.
Engagement rate. Finally, we examine the rate at which participants engaged with the posts they viewed. Unlike the engagement volume analysis, which counts views of at least one second, here we consider all views, since participants could engage even with posts that briefly entered their viewport. We use Beta regression models to appropriately account for the fact that the outcomes are rates. To accommodate rates of zero or one, we transform the outcomes using the commonly used transformation by Smithson and Verkuilen  (?). In addition to the results from the Beta regression models, we also report regression tables from an equivalent analysis using linear models (Tables S14S17). As in the engagement volume analysis, we use a model specification similar to our main analyses (for both the mean and precision parameters) and report average marginal effects.
In the Reduced Exposure experiment, we observe that although participants in the treatment group viewed and liked fewer posts overall, they engage with posts they viewed at higher rates than those in the control condition: 0.61% higher like rate (95% CI: [0.11%, 1.12%]) and 0.058% higher repost rate (95% CI: [0.009%, 0.11%]), as shown in Figure S19 (right panel). In the Increased Exposure experiment, we do not find any significant effects on the like (estimate: 0.22%; 95% CI: [-0.57%, 1.01%]) and repost (estimate: -0.052; 95% CI: [-0.17%, 0.066%]) engagement rates.
Finally, as we discuss in Section S1.3, our experiments were designed to detect meaningful effects on affective polarization and are not necessarily powered to detect small effects on engagement. To contextualize the results, Table S18 reports the observed and post-hoc Minimum Detectable Effects (MDE) for each of the engagement metrics discussed above.
Number of Sessions (Daily)
Reduced Exposure Increased Exposure
(Intercept) $0.43^{*}$ $0.51^{*}$
$[0.32;0.54]$ $[0.37;0.64]$
Condition: Treatment 0.06-0.06 0.04-0.04
[0.16;0.04][-0.16;0.04] [0.16;0.08][-0.16;0.08]
Baseline: Number of Sessions (Daily) $0.68^{*}$ $0.65^{*}$
$[0.63;0.72]$ $[0.59;0.70]$
Recruitment Platform: CloudResearch $0.19^{*}$ $0.22^{*}$
$[0.08;0.29]$ $[0.10;0.34]$
Num. obs. $622$ $468$
Null hypothesis value outside the confidence interval.
Table S6: Engagement: Number of Sessions. Regression tables of the analysis of the number of sessions per day using linear models.
Time Spent (Daily in Minutes)
Reduced Exposure Increased Exposure
(Intercept) $80^{*} 8.8$$89^{*}$
$[5.55;12.06]$ $[4.77;13.01]$
Condition: Treatment $4.71^{*}$ 1.76-1.76
[8.22;1.20][-8.22;-1.20] [6.17;2.66][-6.17;2.66]
Baseline: Time Spent (Daily in Minutes) $0.77^{*}$ $0.79^{*}$
$[0.73;0.81]$ $[0.73;0.84]$
Recruitment Platform: CloudResearch $4.03^{*}$ $3.78$
$[0.39;7.66]$ [0.69;8.25][-0.69;8.25]
Num. obs. $622$ $468$
Null hypothesis value outside the confidence interval.
Table S7: Engagement: Time Spent. Regression tables of the analysis of the time spent in minutes per day using linear models.
Reduced Exposure, Number of Views
Linear Quasi-Poisson
(Intercept) $326.62^{*}$ $6.71^{*}$
$[209.39;443.85]$ $[6.59;6.84]$
Condition: Treatment $213.21^{*}$ $0.15^{*}$
[342.39;84.02][-342.39;-84.02] [0.30;0.01][-0.30;-0.01]
Baseline: Number of Views $1.96^{*}$ $0.00^{*}$
$[1.86;2.07]$ $[0.00;0.00]$
Recruitment Platform: CloudResearch $85.98$ $0.14^{*}$
[47.65;219.61][-47.65;219.61] $[0.00;0.29]$
Num. obs. $622$ $622$
Null hypothesis value outside the confidence interval.
Table S8: Engagement: Reduced Exposure Experiment, Number of Views. Regression results for the number of posts viewed during the 7-day intervention period, using linear and quasi-Poisson models.
Increased Exposure, Number of Views
Linear Quasi-Poisson
(Intercept) $311.00^{*}$ $6.57^{*}$
$[181.04;440.97]$ $[6.44;6.71]$
Condition: Treatment 54.96-54.96 0.13-0.13
[199.62;89.69][-199.62;89.69] [0.28;0.03][-0.28;0.03]
Baseline: Number of Views $1.83^{*}$ $0.00^{*}$
$[1.70;1.96]$ $[0.00;0.00]$
Recruitment Platform: CloudResearch $119.21$ $0.19^{*}$
[27.24;265.66][-27.24;265.66] $[0.04;0.35]$
Num. obs. $468$ $468$
Null hypothesis value outside the confidence interval.
Table S9: Engagement: Increased Exposure Experiment, Number of Views. Regression results for the number of posts viewed during the 7-day intervention period, using linear and quasi-Poisson models.
Reduced Exposure, Number of Likes
Linear Quasi-Poisson
(Intercept) $16.41^{*}$ $3.72^{*}$
$[4.18;28.63]$ $[3.46;3.98]$
Condition: Treatment $19.57^{*}$ $0.47^{*}$
[34.06;5.08][-34.06;-5.08] [0.78;0.16][-0.78;-0.16]
Baseline: Number of Likes $1.93^{*}$ $0.00^{*}$
$[1.85;2.01]$ $[0.00;0.00]$
Recruitment Platform: CloudResearch $9.90$ $0.25$
[5.09;24.88][-5.09;24.88] [0.05;0.56][-0.05;0.56]
Num. obs. $622$ $622$
Null hypothesis value outside the confidence interval.
Table S10: Engagement: Reduced Exposure Experiment, Number of Likes. Regression results for the number of posts liked during the 7-day intervention period, using linear and quasi-Poisson models.
Increased Exposure, Number of Likes
Linear Quasi-Poisson
(Intercept) $2.44$ $3.08^{*}$
[8.17;13.05][-8.17;13.05] $[2.76;3.39]$
Condition: Treatment 0.73-0.73 $0.22$
[13.20;11.74][-13.20;11.74] [0.11;0.55][-0.11;0.55]
Baseline: Number of Likes $2.43^{*}$ $0.01^{*}$
$[2.29;2.56]$ $[0.01;0.01]$
Recruitment Platform: CloudResearch $11.85$ $0.46^{*}$
[0.76;24.47][-0.76;24.47] $[0.14;0.78]$
Num. obs. $468$ $468$
Null hypothesis value outside the confidence interval.
Table S11: Engagement: Increased Exposure Experiment, Number of Likes. Regression results for the number of posts liked during the 7-day intervention period, using linear and quasi-Poisson models.
Reduced Exposure, Number of Reposts
Linear Quasi-Poisson
(Intercept) $4.29^{*}$ $2.13^{*}$
$[1.38;7.21]$ $[1.60;2.66]$
Condition: Treatment 2.45-2.45 0.51-0.51
[5.96;1.07][-5.96;1.07] [1.28;0.25][-1.28;0.25]
Baseline: Number of Reposts $2.07^{*}$ $0.01^{*}$
$[2.02;2.11]$ $[0.00;0.01]$
Recruitment Platform: CloudResearch 1.81-1.81 0.71-0.71
[5.45;1.82][-5.45;1.82] [1.59;0.18][-1.59;0.18]
Num. obs. $622$ $622$
Null hypothesis value outside the confidence interval.
Table S12: Engagement: Reduced Exposure Experiment, Number of Reposts. Regression results for the number of reposts during the 7-day intervention period, using linear and quasi-Poisson models.
Increased Exposure, Number of Reposts
Linear Quasi-Poisson
(Intercept) $0.49$ $0.98^{*}$
[3.27;4.24][-3.27;4.24] $[0.45;1.52]$
Condition: Treatment $0.92$ $0.46$
[3.58;5.41][-3.58;5.41] [0.02;0.93][-0.02;0.93]
Baseline: Number of Reposts $2.27^{*}$ $0.02^{*}$
$[2.14;2.39]$ $[0.02;0.03]$
Recruitment Platform: CloudResearch $0.99$ $0.40$
[3.56;5.55][-3.56;5.55] [0.19;0.99][-0.19;0.99]
Num. obs. $468$ $468$
Null hypothesis value outside the confidence interval.
Table S13: Engagement: Increased Exposure Experiment, Number of Reposts. Regression results for the number of reposts during the 7-day intervention period, using linear and quasi-Poisson models.
Reduced Exposure, Like Rate
Linear Beta
(Intercept) $0.00$ $4.67^{*}$
[0.00;0.01][-0.00;0.01] [4.83;4.50][-4.83;-4.50]
Condition: Treatment $0.00$ $0.22^{*}$
[0.00;0.01][-0.00;0.01] $[0.04;0.40]$
Baseline: Like Rate $0.86^{*}$ $14.66^{*}$
$[0.83;0.89]$ $[13.84;15.48]$
Recruitment Platform: CloudResearch $0.00$ $0.06$
[0.01;0.01][-0.01;0.01] [0.13;0.25][-0.13;0.25]
Precision, (Intercept) $4.08^{*}$
$[3.85;4.30]$
Precision, Condition: Treatment $0.32^{*}$
[0.57;0.07][-0.57;-0.07]
Precision, Baseline: Like Rate $6.91^{*}$
[7.93;5.89][-7.93;-5.89]
Precision, Recruitment Platform: CloudResearch 0.09-0.09
[0.34;0.17][-0.34;0.17]
Num. obs. $622$ $622$
Null hypothesis value outside the confidence interval.
Table S14: Engagement: Reduced Exposure Experiment, Like Rate. Regression results for the like rate, i.e., number of posts liked over number of posts viewed, during the 7-day intervention period, using linear and Beta models.
Increased Exposure, Like Rate
Linear Beta
(Intercept) $0.01$ $4.18^{*}$
[0.00;0.02][-0.00;0.02] [4.38;3.99][-4.38;-3.99]
Condition: Treatment 0.00-0.00 $0.06$
[0.01;0.01][-0.01;0.01] [0.16;0.29][-0.16;0.29]
Baseline: Like Rate $0.82^{*}$ $11.38^{*}$
$[0.76;0.87]$ $[10.44;12.33]$
Recruitment Platform: CloudResearch $0.00$ $0.23$
[0.01;0.01][-0.01;0.01] [0.00;0.46][-0.00;0.46]
Precision, (Intercept) $3.59^{*}$
$[3.33;3.84]$
Precision, Condition: Treatment 0.17-0.17
[0.47;0.12][-0.47;0.12]
Precision, Baseline: Like Rate $5.92^{*}$
[7.05;4.79][-7.05;-4.79]
Precision, Recruitment Platform: CloudResearch $0.36^{*}$
[0.66;0.06][-0.66;-0.06]
Num. obs. $468$ $468$
Null hypothesis value outside the confidence interval.
Table S15: Engagement: Increased Exposure Experiment, Like Rate. Regression results for the like rate, i.e., number of posts liked over number of posts viewed, during the 7-day intervention period, using linear and Beta models.
Reduced Exposure, Repost Rate
Linear Beta
(Intercept) $0.00$ $6.24^{*}$
[0.00;0.00][-0.00;0.00] [6.38;6.11][-6.38;-6.11]
Condition: Treatment $0.00$ $0.19^{*}$
[0.00;0.00][-0.00;0.00] $[0.03;0.35]$
Baseline: Repost Rate $0.98^{*}$ $44.61^{*}$
$[0.95;1.00]$ $[40.38;48.83]$
Recruitment Platform: CloudResearch 0.00-0.00 0.11-0.11
[0.00;0.00][-0.00;0.00] [0.27;0.05][-0.27;0.05]
Precision, (Intercept) $6.22^{*}$
$[6.01;6.43]$
Precision, Condition: Treatment $0.30^{*}$
[0.55;0.05][-0.55;-0.05]
Precision, Baseline: Repost Rate $42.55^{*}$
[46.96;38.14][-46.96;-38.14]
Precision, Recruitment Platform: CloudResearch $0.31^{*}$
$[0.06;0.57]$
Num. obs. $622$ $622$
Null hypothesis value outside the confidence interval.
Table S16: Engagement: Reduced Exposure Experiment, Repost Rate. Regression results for the repost rate, i.e., number of posts reposted over number of posts viewed, during the 7-day intervention period, using linear and Beta models.
Increased Exposure, Repost Rate
Linear Beta
(Intercept) $0.00$ $5.92^{*}$
[0.00;0.00][-0.00;0.00] [6.07;5.78][-6.07;-5.78]
Condition: Treatment $0.00$ 0.08-0.08
[0.00;0.00][-0.00;0.00] [0.28;0.11][-0.28;0.11]
Baseline: Repost Rate $0.89^{*}$ $35.43^{*}$
$[0.81;0.96]$ $[32.04;38.82]$
Recruitment Platform: CloudResearch $0.00$ $0.78^{*}$
[0.00;0.01][-0.00;0.01] $[0.56;1.00]$
Precision, (Intercept) $6.04^{*}$
$[5.80;6.28]$
Precision, Condition: Treatment $0.06$
[0.23;0.35][-0.23;0.35]
Precision, Baseline: Repost Rate $25.04^{*}$
[29.03;21.05][-29.03;-21.05]
Precision, Recruitment Platform: CloudResearch $1.58^{*}$
[1.89;1.28][-1.89;-1.28]
Num. obs. $468$ $468$
Null hypothesis value outside the confidence interval.
Table S17: Engagement: Increased Exposure Experiment, Repost Rate. Regression results for the repost rate, i.e., number of posts reposted over number of posts viewed, during the 7-day intervention period, using linear and Beta models.
Reduced Exposure Increased Exposure
Metric Observed Effect MDE Observed Effect MDE
Number of Sessions (Daily) -0.06 0.14 -0.04 0.17
Time spent (Daily in Minutes) -4.71 5.02 -1.76 6.31
Number of Views (Total) -170.00 249.41 -132.00 254.99
Number of Likes (Total) -24.10 29.84 10.30 27.46
Number of Reposts (Total) -4.18 17.67 3.38 7.33
Like Rate 0.61% 0.79% 0.22% 1.25%
Repost Rate 0.06% 0.08% -0.05% 0.19%
Table S18: Engagement: Observed and post-hoc Minimum Detectable Effects (MDE) in the two experiments (power = 0.8, $\alpha=0.05$). For the engagement volume (number of views, likes, and reports) and engagement rate (like and repost rate) metrics, we first computed the MDE coefficients from the non-linear models used to estimate the treatment effects, and then calculated the corresponding average marginal effects, which are reported here.
S10 Population Average Treatment Effects
To examine the generalizability of our treatment effect estimates, we reweighted our sample to reflect the X user population in terms of party affiliation, education, and race. To determine the proportion of X users across these dimensions, we relied on the 2024 National Public Opinion Reference Survey conducted by Pew Research. The survey was fielded between February 1 and June 10, 2024, just before the start of our experiment (July 7, 2024) and after the rebranding of Twitter to X. Since we did not preregister this analysis, we minimize the number of subjective modeling decisions by considering the same participant characteristics and following a similar reweighting procedure as the Facebook and Instagram Election studies  (?, ?). We reweight our sample based on the following characteristics: (a) Party affiliation: Democrat vs. Republican, considering Independents as partisans depending on which party they lean towards; (b) Education: less than a college degree vs. college degree or more; and (c) Race: white vs. non-white.
We used raking to generate the weights that calibrate to the population estimates of these participant characteristics. We did not trim the weights since there were no extreme values. We consider the participants who completed the post-experiment survey and at least one affective polarization or emotions in-feed survey during the intervention period (N=1,112). Compared to the X population, our sample included fewer Republicans (Reduced Exposure sample: 31.6%, Increased Exposure sample: 33%, X population: 48.4%), fewer participants with less than a college degree (Reduced Exposure sample: 43.7%, Increased Exposure sample: 50.9%, X population: 65.6%), and a similar proportion of white participants (Reduced Exposure sample: 75.7%, Increased Exposure sample: 76.4%, X population: 74.4%). We compute the effective sample size  (?) of the reweighted sample to estimate the proportional increase in variance due to reweighting. We find that our reweighted sample corresponds to a 23.3% smaller sample size in the Reduced Exposure experiment and a 17.3% smaller sample size in the Increased Exposure experiment.
Figure S20: Affective polarization: Comparison between the Population Average Treatment Effect estimates (PATE)—reweighting the sample to be representative of the X population in terms of party affiliation, education, and race—and the Sample Average Treatment Effects (SATE) estimates reported in the main text (Figure 2). The error bars represent 95% confidence intervals. Point estimates for which the 95% confidence intervals include zero are displayed with lighter colors.
Figure S21: Emotions: Comparison between the Population Average Treatment Effect estimates (PATE)—reweighting the sample to be representative of the X population in terms of party affiliation, education, and race—and the Sample Average Treatment Effects (SATE) estimates reported in the main text (Figure 3). The error bars represent 95% confidence intervals. Point estimates for which the 95% confidence intervals include zero are displayed with lighter colors.
To incorporate the raking weights in our analyses, we used the svyglm function from the survey R package  (?, ?). Since we are unaware of any well-established method for incorporating raking weights into mixed-effect models, we first averaged each participant’s responses to the in-feed surveys and then ran weighted regression using svyglm. We note that when we apply the same averaging approach to the unweighted data and run an ordinary linear regression, the resulting point estimates and confidence intervals are similar to those obtained with the mixed-effects models.
We find that the Population Average Treatment Effect (PATE) estimates closely align with the Sample Average Treatment Effect (SATE) estimates (Figures S20 and S21). The significant treatment effects observed for affective polarization and emotional outcomes remain significant after reweighting. The only exception is the treatment effect of reduced exposure to AAPA post-experiment: although the PATE point estimate is similar to the SATE, its confidence interval now includes zero, reflecting a loss of precision from reweighting. In the Increased Exposure experiment, the PATE estimate for affective polarization in-feed is slightly larger than the SATE, while in the Reduced Exposure experiment, it is slightly smaller. The PATE estimates of the treatment effects on emotions are similar to the SATE estimates.
S11 Supplementary Results
S11.1 Affective Polarization: Regression Tables
@@
The preregistered regression analysis for the in-feed effect is reported in Table S19, and the regression analyses for the post-experiment effect are reported in Table S20.
@@
Table S21 and Table S22 show the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-experiment survey or only at least one in-feed survey. For both experiments, the treatment effect remains statistically significant.
We note that in this sample there is also no evidence of covariate imbalance (reduce, post-experiment: $p=0.86$; reduce, in-feed: $p=0.59$; increase, post-experiment: $p=0.77$; increase, in-feed: $p=0.31$). The attrition rates were as follows—reduce, post-experiment, treatment: 27.7% vs. control: 26.1%; reduce, in-feed, treatment: 33.1% vs. control: 36.4%; increase, post-experiment, treatment: 19.4% vs. control: 22.0%; increase, in-feed, treatment: 26.2% vs. control: 24.2%. We also find no evidence of differential attrition rates (reduce, post-experiment: $p=0.55$; reduce, in-feed: $p=0.28$; increase, post-experiment: $p=0.4$; increase, in-feed: $p=0.56$), or differential attrition patterns (reduce, post-experiment: $p=0.49$; reduce, in-feed: $p=0.69$; increase, post-experiment: $p=0.78$; increase, in-feed: $p=0.98$). The testing procedures are described in Section S1.3.
@@
Recruitment Platform: CloudResearch 0.58-0.58 1.58-1.58
[2.69;1.54][-2.69;1.54] [3.58;0.43][-3.58;0.43]
@@
Table S19: Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
@@
Recruitment Platform: CloudResearch 0.76-0.76 $2.72^{*}$
[2.79;1.28][-2.79;1.28] [5.06;0.37][-5.06;-0.37]
@@
Table S20: Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
@@
Recruitment Platform: CloudResearch 0.58-0.58 0.82-0.82
[2.68;1.51][-2.68;1.51] [2.84;1.19][-2.84;1.19]
@@
Table S21: Full population – Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
@@
Recruitment Platform: CloudResearch 0.79-0.79 $3.07^{*}$
[2.72;1.14][-2.72;1.14] [5.30;0.84][-5.30;-0.84]
@@
Table S22: Full population – Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
S11.2 Emotions: Regression Tables
The preregistered regression analysis for the in-feed effect in the reduced exposure is reported in Table S23 and increased exposure in Table S24. The post-experiment effect regression results for the reduced exposure experiment are reported in Table S25 and for the increased exposure in Table S26.
Table S27, Table S28, Table S29, and Table S30 report the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-experiment survey or only at least one in-feed survey. The results are qualitatively similar, the variation in negative emotions is statistically significant in both experiments: increased in the increased exposure experiment and reduced in the reduced exposure experiment.
@@
Recruitment Platform: CloudResearch $0.56$ $2.42$ $5.05^{*}$ 3.09-3.09
[2.16;3.27][-2.16;3.27] [0.41;5.24][-0.41;5.24] [8.54;1.55][-8.54;-1.55] [6.25;0.08][-6.25;0.08]